A carrier signal c(?) = ? cos(??) is modulated by a single-tone signal ?(?) = ? ???(??) to form the AM signal: s(t)= [? + ????(??)]???(??), where ? = 10, ? = 2, ? = 1000 ??, ? = 50 ??. a- Obtain an expression for the sidebands of ?(?). b- Sketch ?(?) c- Sketch the spectra |?(?)|, |?(?)|, |?(?)| against f. d- Calculate the modulation index (m) e- What bandwidth is required to transmit the AM signal. f- Show that the sideband power to the power in the modulated wave is equal to   
Answer #1

Giuon corrion signal ccé) - A cont.) whore = A caug41) A = 10 t = 1000 Hz Then Cee) = 10 collar tegut) Similanly me) = Acout)CQS (A+B) + las (4-3) las SC) = 10 [1+0.2 les art so t)] cas (att imot) 10 cos art 1 rook + 2 ces (artist) LOJ 211 sot kle kn(6) Sketch of sct) Envelope el wyt on (72S 8 AC-Am Ac=10 Am=2Francier Transforms of Owot e mct)=cas Wort 2. cas Wol is - Twol f e 2 2 f[emaet J - 21 600-ws) = 77 80w-wo) So f [ casuot २MGH), 501-6) +504 +40) 2 Speettern of Mcd) es Almeldi 71 (te) given mct) = a las 2150 t) Then MCA)=2864-50) + 2864 +50) 2 2 5Samilarly SCL) = 10 cos Cartierot) + cos car 1000 €) + cas (21 950 €) SCH) = 5814-16) +55(++lroo) t 567-1050) +55(++1050) + *(d) Modulation inden 0.2 , as per the general teten af Am signal . Modulation inden (m) = Am ale Ac حياة 10 0.2 ce) Bandouldt